OpenStudy (anonymous):

There are 9red,7black and 5 white balls in an urn.Three balls are taken at random from it, one after another without replacement . find the probability that they are red,black and white respectively.

OpenStudy (klimenkov):

What is the probability to take red ball from this urn?

OpenStudy (anonymous):

@avivora the probability that they are red,black and white respectively is C(9,1)/C(21,1) * C(7,1)/C(20,1) *C(5,1)/C(19,1) = 9/21 * 7/20 * 5/19=3/76

OpenStudy (anonymous):

RIGHT ANS 3/76